OpenStudy (anonymous):

Under ideal conditions, the number of ecoli bacteria can double every 20 minutes. This behavior can be modeled by the exponential function N(t)=Nzero(2^0.05t) where t is the time in minutes andNzero is the initial number of bacteria. A) if the initial number of bacteria is 2, how many will be present in 4 hours?

OpenStudy (amistre64):

N{0} = 2; and t=4, we can googlize the equation and itll pop out an answer :)

OpenStudy (amistre64):

2(2^(.05*4))= 2.29739671 http://www.google.com/search?aq=f&sourceid=chrome&ie=UTF-8&q=2(2^(.05*4))

OpenStudy (amistre64):

might have to present t in minutes tho right?

OpenStudy (amistre64):

3 sets of 20min per hour; 4 hours then equals 12 cycles

OpenStudy (anonymous):

do you change the time to minutes, or leave it in hours? because it is telling me the correct answer is 8,192

OpenStudy (anonymous):

we were typing at the same time :-)

OpenStudy (amistre64):

i believe you change it to minutes

OpenStudy (amistre64):

let me see if I can work up a better looking formula ; or at least verify theirs

OpenStudy (amistre64):

4 = 2e^(k*20) ln(4/2) ------ = k 20

OpenStudy (anonymous):

aha! thank you.

OpenStudy (amistre64):

i get a rate of .03466

OpenStudy (amistre64):

but im gonna brute math it :) a{t} = amount a{0:00} = 2 a{0:20} = 4 a{0:40} = 8 a{1:00} = 16 a{1:20} = 32 a{1:40} = 64 a{2:00} = 128 a{2:20} = 256 a{2:40} = 512 a{3:00} = 1024 a{3:20} = 2048 a{3:40} = 4096 a{4:00} = 8192

OpenStudy (amistre64):

A = 2e^((ln(2)/20)*240) = 8,191.99999....